how do i solve x+y=46
1
Expert's answer
2011-12-01T11:09:04-0500
Here you can get y = 46 - x. The solution of this equation is the whole line.
For example if we pick x=1 then y=46-1=45
if x=1/3 then y=46-1/3=45 2/3